Being F the subset of M3x3(R)of the hemi-symmetric matrices ( such as A=-AT).

i) Show that F is a subspace of M3x3(R).

ii) Determine the dimension of F.

iii) Determine the base of F.

iv) Being T:F-R3 the application that corresponds to each matrix A= (aij) of  F the vector (+12+ +13, +18 + 41, 42) of .

Determine the matrix that represents T regarding the base of the previous question (iii) and the canonical base of .

v) Determine if T is injective.

vi) Determine if T is surjective.

vii) Verify that T satisfies the Dimension theorem , such as dim F = dim Nuc T+dim Im T
Answer #1

F = {A E M3x3. (IR) A = -AT} . (1) We just need to show closure. Let A BE F Ee - DER then (A+ dB) = A + 2Br = -A-4B. - - (A+Q(vi) Let (x, y, z) ER² then (x, y, z) = xe, tye, + Zes Also T(E)-T(E)= .. & T(E)- (T(E)-T(E)) = e. so xT(E) tag (T(E₂) -T(E)+please rate